# Semiclassical measure of the propagation between two topological insulators

### v2

@@ Appendix A Proof of compactness at infinity
Let $T>0$. If $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is compact at infinity then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to (1) is compact at infinity uniformly for $t\in[-T,T]$.
@@
Appendix B Construction of semiclassical measures
B.1 Semiclassical measures for families valued in a separable Hilbert space
We will use the more general framework of [14] and consider families valued in a separable Hilbert space $\mathcal{H}$. More precisely, in this section, $\mathcal{H}=\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$. Note that in section 2.1, $\mathcal{H}=\mathbb{C}^{2}$.
Lemma B.1 (Operator valued measure).
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ a bounded family in $\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)$, then there exists ${(\varepsilon_{k})_{k>0}}$ a vanishing sequence and a positive operator valued measure $\Gamma\mathrm{d}\gamma$ such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$,
$\left\langle\mathrm{Op}_{\varepsilon_{k}}(a)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}a(x,\xi)\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi),$
@@ B.1 Semiclassical measures for families valued in a separable Hilbert space
$\sum_{n\in\mathbb{N}}\left\lVert f^{\varepsilon}_{n}\right\rVert_{\mathrm{L}^{2}(\mathbb{R})}^{2}=\left\lVert f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)}^{2}\leqslant C.$ (36)
@@
First of all, we construct the operator valued measure $\Gamma\mathrm{d}\gamma$. By (36),
@@
then ${\mathrm{d}\gamma\coloneqq\sum_{n\in\mathbb{N}}\mathrm{d}\mu_{n,n}}$ defines a positive Radon measure of finite mass. By the Radon-Nykodym Theorem, there exists a sequence of $\mathrm{d}\gamma$-measurable functions $\left(\Gamma_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$ such that $\forall(n,m)\in\mathbb{N}^{2}$,
$\mathrm{d}\mu_{n,m}=\Gamma_{n,m}\mathrm{d}\gamma$
with $0\leqslant\Gamma_{n,n}\leqslant 1$ and $\Gamma_{n,m}=\overline{\Gamma_{m,n}}$ $\mathrm{d}\gamma$-almost everywhere. Let us define the operator valued function $\Gamma:(x,\xi)\mapsto\Gamma(x,\xi)$ such that, for all $(n,m)\in\mathbb{N}^{2}$, for all $(x,\xi)\in\mathbb{R}^{2}$,
$\left\langle\Gamma(x,\xi)h_{n},h_{m}\right\rangle_{\mathcal{H}}=\Gamma_{n,m}(x,\xi).$
Because of the positivity of $\left(\mathrm{d}\mu_{n,n}\right)_{n\in\mathbb{N}}$ and the adjoint relation of $\left(\mathrm{d}\mu_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$, $\Gamma(x,\xi)$ is a positive self-adjoint operator and its trace satisfies
$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi)=\sum_{n\in\mathbb{N}}\int_{\mathbb{R}^{2}}\Gamma_{n,n}(x,\xi)\mathrm{d}\gamma(x,\xi)=\sum_{n\in\mathbb{N}}\mathrm{d}\gamma_{n,n}\left(\mathbb{R}^{2}\right)<+\infty.$
Therefore, the pair $\left(\Gamma,\mathrm{d}\gamma\right)$ generates an operator measure $\Gamma\mathrm{d}\gamma$. We now have to prove that $\Gamma\mathrm{d}\gamma$ plays the expected role.
@@
$\sup_{(x,\xi)\in\mathrm{supp}(a)}\left\lVert a(x,\xi)-a_{n}(x,\xi)\right\rVert_{\mathcal{L}\left(\mathcal{H}\right)}\underset{n\to+\infty}{\longrightarrow}0.$ (37)
@@
$\underset{k\to+\infty}{\longrightarrow}\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}(x,\xi)h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\mathrm{d}\gamma_{m,\ell}(x,\xi)$
@@
$=\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\Gamma_{m,\ell}\mathrm{d}\gamma(x,\xi)$
$=\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}a_{n}(x,\xi)\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi).$
Then, with (37) and Lebesgue dominated convergence Theorem, we have
$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big{(}\big{(}a(x,\xi)-a_{n}(x,\xi)\big{)}\Gamma(x,\xi)\Big{)}\mathrm{d}\gamma(x,\xi)\underset{n\to+\infty}{\longrightarrow}0.$
B.2 Two-scale semiclassical measures
In this section, we prove Theorem 3.3.
Proof of Theorem 3.3.
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ uniformly bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. For simplicity, we assume that the sequence $\left(\varepsilon\right)_{\varepsilon>0}$ realizes the semiclassical measure $\mathrm{d}\mu$ of ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$. We use the decomposition (17). First we analysis of the concentration of $a^{\delta}$.
  • $\bullet$
    Construction of $\mathrm{d}\mu$ outside $\mathcal{C}$
For $a\in\mathfrak{A}$ such that $a=0$ in a neighborhood of $\mathcal{C}$, then for $\varepsilon$ small enough, we have
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\right)$ $=\mathrm{Op}_{\varepsilon}\left(a\left(s,y;\sigma,\eta;\frac{y}{\sqrt{\varepsilon}},\frac{\sigma}{\sqrt{\varepsilon}},\frac{\eta}{\sqrt{\varepsilon}}\right)\right)$
$=\mathrm{Op}_{1}\left(a_{\infty}\left(s,y,\varepsilon\sigma,\varepsilon\eta,\frac{(y,\varepsilon\sigma,\varepsilon\eta)}{|(y,\varepsilon\sigma,\varepsilon\eta)|}\right)\right).$
Then we construct the pair $\Gamma\mathrm{d}\gamma$.
  • $\bullet$
    Construction of $\Gamma\mathrm{d}\gamma$
Let $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{7},\mathbb{C}^{2,2}\right)$, since $a$ is compactly supported, we recall (21)
$\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a\right)=\mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{\sqrt{\varepsilon},1}(b)\mathfrak{T}_{\sqrt{\varepsilon}}+\mathcal{O}\left(\sqrt{\varepsilon}\right),$
The function $b$ is valued in the set of compact operators on $\mathcal{H}$. We can see ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ as a bounded family of $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$. We are thus doing classical calculus on $\mathrm{L}^{2}\left(\mathbb{R}_{s},\mathcal{H}\right)$. We apply Lemma B.1, then there exists $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$ a vanishing sequence such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{7},\mathbb{C}^{2,2}\right)$,
$\lim_{k\to+\infty}\left\langle\mathfrak{Op}_{\varepsilon_{k}}\left(a\right)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})}=\int_{\mathbb{R}^{2}}\mathrm{Tr}\Big{(}a^{W}(s,0;0,0;y,\sigma,\mathrm{D}_{y})\Gamma(s,\sigma)\Big{)}\mathrm{d}\gamma(s,\sigma).$
Then, we apply this property to $a_{\varepsilon,R}$. Finally we construct the pair $M_{\infty}\mathrm{d}\nu_{\infty}$.
  • $\bullet$
    Construction of $M_{\infty}\mathrm{d}\nu_{\infty}$
Let $a\in\mathfrak{A}$, then we recall (20)
$\begin{split}\mathfrak{Op}^{\mathcal{C}}_{\varepsilon}\left(a_{\varepsilon,\delta}^{R}\right)=&\ \mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\Bigg{(}a(s,\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta)|}{\delta}\right)\\ &\times\left(1-\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\Bigg{)}\mathfrak{T}_{\sqrt{\varepsilon}}\\ =&\ \mathfrak{T}_{\sqrt{\varepsilon}}^{\dagger}\mathrm{Op}_{1}\underbrace{\left(a(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\chi\left(\frac{|(\sqrt{\varepsilon}y,\varepsilon\sigma,\sqrt{\varepsilon}\eta)|}{\delta}\right)\left(1-\chi\left(\frac{|(y,\sqrt{\varepsilon}\sigma,\eta)|}{R}\right)\right)\right)}_{\eqqcolon\widetilde{a}^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)}\mathfrak{T}_{\sqrt{\varepsilon}}\\ &+\mathcal{O}\left(\delta\right).\end{split}$
therefore we consider the family $\widetilde{f}^{\varepsilon}\coloneqq\mathfrak{T}_{\sqrt{\varepsilon}}f^{\varepsilon}$ which is bounded in $\mathrm{L}^{2}\left(\mathbb{R}^{2},\mathbb{C}^{2}\right)$. Then the quantity
is uniformly bounded in $\varepsilon>0$ and $R>1$. $a^{R}_{\delta}$ is supported at distance at least $R\sqrt{\varepsilon}$ of $\mathcal{C}$, so
$\mathscr{I}^{\varepsilon_{k}}_{\widetilde{f}}\left(\widetilde{a}^{R}_{\delta}\right)=\left\langle\mathrm{Op}_{1}\left(a_{\infty}(s,0,0,0,y,\sqrt{\varepsilon_{k}}\sigma,\eta)\right)\widetilde{f}^{\varepsilon_{k}},\widetilde{f}^{\varepsilon_{k}}\right\rangle_{\mathrm{L}^{2}(\mathbb{R}^{2},\mathbb{C}^{2})},$
as soon as $R$ is large enough, depending of $\varepsilon$. We then deduce by a diagonal extraction argument that there exists a subsequence $\left(\varepsilon_{n}\right)_{n\in\mathbb{N}}$ of $\left(\varepsilon_{k}\right)_{k\in\mathbb{N}}$, a sequence $\left(R_{n}\right)_{n\in\mathbb{N}}$ and a sequence $\left(\delta_{n}\right)_{n\in\mathbb{N}}$ such that for all $a\in\mathfrak{A}$,
It remains to prove that $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a measure, which will allow us to define $M_{\infty}\mathrm{d}\nu_{\infty}$. First, we prove that $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a non negative matrix-valued distribution. We observe that there exists $C>0$ such that for all $a\in\mathfrak{A}$, for all $n\in\mathbb{N}$,
with $\mathrm{N}_{d}$ a Schwarz semi-norm. We have $\mathrm{N}_{d}\left(\widetilde{a}^{R_{n}}_{\delta_{n}}\right)\to\mathrm{N}_{d}\left(a_{\infty}\right)$ then $\mathscr{I}\left(a_{\infty}\right)\leqslant C\mathrm{N}_{d}\left(a_{\infty}\right)$. Therefore $\mathscr{I}$ is a matrix-valued distribution. Secondly, we observe that the operators $a\mapsto\mathrm{Op}_{1}\left(a^{R}_{\delta}(s,0,0,0,y,\sqrt{\varepsilon}\sigma,\eta)\right)$ satisfy a semiclassical calculus in the parameters $\varepsilon$, $1/R$ and $\delta$. We have for all $a,a_{1},a_{2}\in\mathfrak{A}$,
  • $\bullet$
  • $\bullet$
  • $\bullet$
Therefore $a_{\infty}\mapsto\mathscr{I}\left(a_{\infty}\right)$ is a non negative matrix-valued measure defined on ×$\mathbb{R}\times\mathbb{S}^{2}$ that we denote $M\mathrm{d}\nu$. ∎
Appendix C Spectral Analysis of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$
In this appendix, we study the spectrum of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ introduced in (34).
Theorem C.1 (Spectrum of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$).
For all $(s,\sigma)\in\mathbb{R}^{2}$, the spectrum of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ consists in multiplicity one eigenvalues
@@ Appendix C Spectral Analysis of the operator $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$
$\mathrm{Sp}\Big{(}\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)\Big{)}=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\},$
with for all $n\in\mathbb{Z}_{\neq 0}$,
$\lambda_{0}(\sigma)\coloneqq-\sigma\;,\;\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n|r(s)^{2}}.$
@@
Moreover, an Hilbertian basis of $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$, composed of eigenfunctions of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$, is given by ${\left(g_{n}^{s,\sigma}\right)_{n\in\mathbb{Z}}}$
$g_{0}^{s}(y)$ $\coloneqq\frac{1}{\sqrt{r(s)}}\begin{pmatrix}0\\ \mathfrak{h}_{0}\big{(}r(s)y\big{)}\end{pmatrix},$
@@
$g_{n}^{s,\sigma}(y)$ $\coloneqq\alpha_{n}(s,\sigma)\begin{pmatrix}\mathfrak{h}_{|n|-1}\big{(}r(s)y\big{)}\\ \frac{\sqrt{2|n|}r(s)}{\lambda_{n}(\sigma)+\sigma}\mathfrak{h}_{|n|}\big{(}r(s)y\big{)}\end{pmatrix},\quad\quad n\in\mathbb{Z}_{\neq 0}$
@@
$\mbox{with}\quad\quad\alpha_{n}(s,\sigma)$ $\coloneqq\frac{1}{\sqrt{2r(s)}}\sqrt{1+\frac{\sigma}{\lambda_{n}(s,\sigma)}},$
@@
and $\left(\mathfrak{h}_{n}\right)_{n\in\mathbb{N}}$ are Hilbert functions defined in (27). Moreover, the projectors over the eigenspaces of $\mathrm{T}_{\mathrm{E}}^{\rm ds}(s,\sigma)$ defined by

### v3

@@ A.1 Application to Dirac equation
Let $T>0$. If $\left(\psi^{\varepsilon}_{0}\right)_{\varepsilon>0}$ is compact at infinity then $\left(\psi^{\varepsilon}_{t}\right)_{\varepsilon>0}$ solution to (1.1) is compact at infinity uniformly for $t\in[-T,T]$.
@@
$\mathrm{tr}\int_{\mathbb{R}^{4}}\mathrm{d}\mu(x,\xi)=1.$
A.2 Semiclassical measures for families valued in a separable Hilbert space
We will use the more general framework of Gérard (1991) and consider families valued in a separable Hilbert space $\mathcal{H}$. More precisely, in this section, $\mathcal{H}=\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$. Note that in Section 1.2, $\mathcal{H}=\mathbb{C}^{2}$.
Lemma A.4 (Operator valued measure).
Let ${\left(f^{\varepsilon}\right)_{\varepsilon>0}}$ a bounded family in $\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)$, then there exists a vanishing sequence of positive numbers ${(\varepsilon_{k})_{k>0}}$ and a positive operator valued measure $M\mathrm{d}\nu$ such that for all $a\in\mathscr{C}_{c}^{\infty}\left(\mathbb{R}^{2},\mathcal{K}\left(\mathcal{H}\right)\right)$,
$\left\langle\mathrm{Op}_{\varepsilon_{k}}(a)f^{\varepsilon_{k}},f^{\varepsilon_{k}}\right\rangle\underset{k\to+\infty}{\longrightarrow}\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big(a(x,\xi)M(x,\xi)\Big)\mathrm{d}\nu(x,\xi),$
@@ A.2 Semiclassical measures for families valued in a separable Hilbert space
$\sum_{n\in\mathbb{N}}\left\lVert f^{\varepsilon}_{n}\right\rVert_{\mathrm{L}^{2}(\mathbb{R})}^{2}=\left\lVert f^{\varepsilon}\right\rVert_{\mathrm{L}^{2}\left(\mathbb{R},\mathcal{H}\right)}^{2}\leqslant C.$ (A.1)
@@
First of all, we construct the operator valued measure $M\mathrm{d}\nu$. By (A.1),
@@
then ${\nu\coloneqq\sum_{n\in\mathbb{N}}\mu_{n,n}}$ defines a positive Radon measure of finite mass. By the Radon-Nykodym Theorem, there exists a sequence of $\mathrm{d}\nu$-measurable functions $\left(M_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$ such that $\forall(n,m)\in\mathbb{N}^{2}$,
$\mathrm{d}\mu_{n,m}=M_{n,m}\mathrm{d}\nu$
with $0\leqslant M_{n,n}\leqslant 1$ and $M_{n,m}=\overline{M_{m,n}}$ $\mathrm{d}\nu$-almost everywhere. Let us define the operator valued function $M:(x,\xi)\mapsto M(x,\xi)$ such that, for all $(n,m)\in\mathbb{N}^{2}$, for all $(x,\xi)\in\mathbb{R}^{2}$,
$\left\langle M(x,\xi)h_{n},h_{m}\right\rangle_{\mathcal{H}}=M_{n,m}(x,\xi).$
Because of the positivity of $\left(\mathrm{d}\mu_{n,n}\right)_{n\in\mathbb{N}}$ and the adjoint relation of $\left(\mathrm{d}\mu_{n,m}\right)_{(n,m)\in\mathbb{N}^{2}}$, $M(x,\xi)$ is a positive self-adjoint operator and its trace satisfies
$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big(M(x,\xi)\Big)\mathrm{d}\nu(x,\xi)=\sum_{n\in\mathbb{N}}\int_{\mathbb{R}^{2}}M_{n,n}(x,\xi)\mathrm{d}\nu(x,\xi)=\sum_{n\in\mathbb{N}}\mathrm{d}\nu_{n,n}\left(\mathbb{R}^{2}\right)<+\infty.$
Therefore, the pair $\left(M,\mathrm{d}\nu\right)$ generates an operator measure $M\mathrm{d}\nu$. We now have to prove that $M\mathrm{d}\nu$ plays the expected role.
@@
$\sup_{(x,\xi)\in\mathrm{supp}(a)}\left\lVert a(x,\xi)-a_{n}(x,\xi)\right\rVert_{\mathcal{L}\left(\mathcal{H}\right)}\underset{n\to+\infty}{\longrightarrow}0.$ (A.2)
@@
$\underset{k\to+\infty}{\longrightarrow}\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}(x,\xi)h_{m},h_{\ell}\right\rangle_{\mathcal{H}}\mathrm{d}\nu_{m,\ell}(x,\xi)$
@@
$=\sum_{m,\ell\leqslant n}\int_{\mathbb{R}^{2}}\left\langle a_{n}h_{m},h_{\ell}\right\rangle_{\mathcal{H}}M_{m,\ell}\mathrm{d}\nu(x,\xi)$
$=\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big(a_{n}(x,\xi)M(x,\xi)\Big)\mathrm{d}\nu(x,\xi).$
Then, with (A.2) and Lebesgue dominated convergence Theorem, we have
$\int_{\mathbb{R}^{2}}\mathrm{Tr}_{\mathcal{L}(\mathcal{H})}\Big(\big(a(x,\xi)-a_{n}(x,\xi)\big)M(x,\xi)\Big)\mathrm{d}\nu(x,\xi)\underset{n\to+\infty}{\longrightarrow}0.$
Appendix B Spectral Analysis of the principal normal form operator
In this appendix, we study the operator $\mathrm{T}_{\mathrm{E}}(s,\sigma)$ introduced in (1.12). First, we rescale the operator by $r(s)$. This is possible because $r$ is smooth ($\mathrm{m}$ is in $\mathscr{C}^{\infty}\left(\mathbb{R}^{2},\mathbb{R}\right)$) and strictly non-negative (Assumption 1.1).
Let us define
$\mathscr{T}\coloneqq\begin{pmatrix}\zeta&\mathfrak{a}^{\dagger}\\ \mathfrak{a}&-\zeta\end{pmatrix}$,
where $\mathfrak{a}\coloneqq y+\mathrm{i}\mathrm{D}_{y}$ denotes the usual annihilation operator on the variable $y$.
Remark B.1.
@@ Appendix B Spectral Analysis of the principal normal form operator
By denoting by $\zeta$ the ratio $\sigma/r(s)$, we have the following, for all $f\in\mathrm{L}^{2}\left(\mathbb{R},\mathbb{C}^{2}\right)$,
$\mathrm{T}_{\mathrm{E}}(s,\sigma)\Bigg[f\Big(r(s)y\Big)\Bigg]$ $=\left[\begin{pmatrix}\sigma&r(s)(y-\mathrm{i}\mathrm{D}_{y})\\ r(s)(y+\mathrm{i}\mathrm{D}_{y})&-\sigma\end{pmatrix}f\right]\Big(r(s)y\Big)$
$=r(s)\left[\begin{pmatrix}\zeta&y-\mathrm{i}\mathrm{D}_{y}\\ y+\mathrm{i}\mathrm{D}_{y}&-\zeta\end{pmatrix}f\right]\Big(r(s)y\Big).$
Then the link between spectra is
$\mathrm{Sp}\Big(\mathrm{T}_{\mathrm{E}}(s,\sigma)\Big)=r(s)\mathrm{Sp}(\mathscr{T}).$
So we can study $\mathscr{T}$ instead of $\mathrm{T}_{\mathrm{E}}(s,\sigma)$.
Let us recall usual properties of creation and annihilation operators.
Proposition B.2.
We have the following.
  • $\mathfrak{a}\mathfrak{a}^{\dagger}=-\partial_{x}^{2}+x^{2}+1$.
  • $\operatorname{sp}\left(\mathfrak{a}\mathfrak{a}^{\dagger}\right)=\{2n\mid n\in\mathbb{N}_{>0}\}$.
  • The Hermite functions $(\mathfrak{h}_{n})_{n\in\mathbb{N}}$, defined in (1.13), form an Hilbertian basis of $\mathrm{L}^{2}(\mathbb{R},\mathbb{C})$, of eigenfunctions of $\mathfrak{a}\mathfrak{a}^{\dagger}$, and we have
    $\mathfrak{a}\mathfrak{a}^{\dagger}\mathfrak{h}_{n}=2(n+1)\mathfrak{h}_{n},\quad n\in\mathbb{N}.$
  • Moreover $\mathfrak{a}$ and $\mathfrak{a}^{\dagger}$ satisfies the following identities
    $\mathfrak{a}\mathfrak{h}_{n}=\sqrt{2n}\mathfrak{h}_{n-1},\quad\mathfrak{a}^{\dagger}\mathfrak{h}_{n}=\sqrt{2(n+1)}\mathfrak{h}_{n+1}.$
The next Proposition consists to diagonalize, at fixed $(s,\sigma)$, the operator $\mathrm{T}_{\mathrm{E}}(s,\sigma)$ on $\mathrm{L}^{2}(\mathbb{R}_{y},\mathbb{C}^{2})$ since $(\mathrm{T}_{\mathrm{E}}(s,\sigma),\mathcal{B}^{1}(\mathbb{R}_{y},\mathbb{C}^{2}))$ is self-adjoint with compact resolvent.
Theorem B.3 (Spectrum of $\mathrm{T}_{\mathrm{E}}(s,\sigma)$).
The spectrum of $\mathrm{T}_{\mathrm{E}}(s,\sigma)$ consists in multiplicity one eigenvalues
$\mathrm{Sp}\Big(\mathrm{T}_{\mathrm{E}}(s,\sigma)\Big)=\left\{\lambda_{0}(\sigma)\right\}\bigcup\left\{\lambda_{n}(s,\sigma)\ |\ n\in\mathbb{Z}_{\neq 0}\right\},\quad(s,\sigma)\in\mathbb{R}^{2},$
with, for all $n\in\mathbb{Z}_{\neq 0}$,
$\lambda_{0}(\sigma)\coloneqq\sigma,\quad\lambda_{n}(s,\sigma)\coloneqq\mathrm{sgn}(n)\sqrt{\sigma^{2}+2|n|r(s)^{2}}.$
@@
Moreover, an Hilbertian basis of $\mathrm{L}^{2}\left(\mathbb{R}_{y},\mathbb{C}^{2}\right)$, composed of eigenfunctions of $\mathrm{T}_{\mathrm{E}}(s,\sigma)$, is given by ${\left(g_{n}^{s,\sigma}\right)_{n\in\mathbb{Z}}}$
$g_{0}^{s}(y)$ $\coloneqq\frac{1}{\sqrt{r(s)}}\begin{pmatrix}\mathfrak{h}_{0}\big(r(s)y\big)\\ 0\end{pmatrix}$,
@@
$g_{n}^{s,\sigma}(y)$ $\coloneqq\alpha_{n}(s,\sigma)\begin{pmatrix}\frac{\sqrt{2|n|}r(s)}{\lambda_{n}(\sigma)-\sigma}\mathfrak{h}_{|n|}\big(r(s)y\big)\\ \mathfrak{h}_{|n|-1}\big(r(s)y\big)\end{pmatrix},\quad\quad n\in\mathbb{Z}_{\neq 0}$
@@
$\mbox{with}\quad\quad\alpha_{n}(s,\sigma)$ $\coloneqq\frac{1}{\sqrt{2r(s)}}\sqrt{1-\frac{\sigma}{\lambda_{n}(s,\sigma)}},$
@@
and $\left(\mathfrak{h}_{n}\right)_{n\in\mathbb{N}}$ are Hermite functions defined in (1.13). Moreover, the projectors over the eigenspaces of $\mathrm{T}_{\mathrm{E}}(s,\sigma)$ defined by